# Biosecurity Safeguards for Generative AI

### v4

@@ Abstract
Additionally, we identify the need for normalizing the unlearning gradients, and also confirm the usefulness of meta-learning. We combine these insights into MUDMAN (Meta-Unlearning with Disruption Masking and Normalization) and validate its effectiveness at preventing the recovery of dangerous capabilities. MUDMAN outperforms the prior TAR method by 40%, setting a new state-of-the-art for robust unlearning. Code: anonymous.4open.science/r/MUDMAN
@@ 1 Introduction
Language models can acquire dangerous skills during pre-training, such as manipulation, hacking and even knowledge useful for creating bioweapons (Li et al., 2024). They may also learn about the safeguards used to control them, which in the future may enable them to subvert such safeguards (Greenblatt et al., 2024; Roger, 2024).
Studies show that popular safety fine-tuning techniques like DPO and RLHF fail to remove dangerous knowledge; instead, they minimally modify model weights to hide it (Lee et al., 2024), allowing it to reemerge through jailbreak inputs (Zou et al., 2023), or even accidentally (Qi et al., 2023; Deeb and Roger, 2024). Even specialized unlearning techniques turn out to be easily reversible (Lynch et al., 2024; Łucki et al., 2025; Deeb and Roger, 2024). It remains unclear which components of unlearning algorithms can make harmful capabilities truly irreversible.
@@ 2 Related work
Some unlearning approaches disrupt intermediate activations within the model (Zou et al., 2024; Li et al., 2024; Rosati et al., 2024), while others attempt to locate and ablate weights responsible for unwanted behavior Wang et al. (2024); Wu et al. (2023); Uppaal et al. (2024); Suau et al. (2024). However, Lo et al. (2024) found that even when unwanted concepts are directly removed, the model can quickly learn to represent them again using neurons with similar meaning.
To address this, modern unlearning techniques increasingly incorporate MAML (model-agnostic meta-learning) (Tamirisa et al., 2024a; Henderson et al., 2023; Tamirisa et al., 2024b). This approach anticipates how an attacker could relearn the target capability, by deriving unlearning gradients from a copy of the model trained on the forget set (Finn et al., 2017).
However, for each existing unlearning methods, there are still ways to elicit the supposedly removed capabilities, such as jailbreak prompts, few-shot prompting, fine-tuning, in-context learning, out-of-distribution inputs or disabling refusal mechanisms with representation engineering (Lynch et al., 2024; Łucki et al., 2025).
@@ 3 Experiment setup
We conduct experiments with three language models of different sizes: pythia-14m (Biderman et al., 2023). SmolLM-135M (Allal et al., 2024) and Llama-3.2-1B (Grattafiori et al., 2024). Using pythia-14m allowed us to iterate fast on our techniques, then we validate our findings on the larger, more modern models.
@@
We demonstrate skill unlearning by attempting to unlearn Python coding ability by using function examples from CodeSearchNet (Husain et al., 2019), while retaining the performance on Wikitext (Merity et al., 2016). To ensure we do not unintentionally unlearn English, we remove comments and docstrings from the Python dataset. While programming is not a harmful behavior, using loss on the Python dataset provides a high signal-to-noise ratio for comparing unlearning methods and serves as a good baseline.
Then we move to knowledge unlearning with a realistic harmful target: unlearning biohazardous knowledge using Pile-Bio (Tamirisa et al., 2024a), a subset of the Pile (Gao et al., 2020) containing texts about molecular biology. The rest of the Pile serves as the retain set.
To test unlearning effectiveness on Pile-Bio, we measure the accuracy on WMDP-Bio (Li et al., 2024), a dataset of 1273 multiple-choice questions designed as a proxy measurement of hazardous biosecurity knowledge. We use temperature=0 to generate the answers.
@@
Method performance is highly dependent on hyperparameter selection. To ensure a fair comparison, we perform an automated hyperparameter search for each tested method using Optuna (Akiba et al., 2019). We verify that each search converges and that the chosen hyperparameter ranges are not saturated. Each search includes hundreds of trials, with each trial consisting of an unlearning stage followed by a fixed relearning stage, and the amount of compute in each stage is constant. (See Appendix G for details.)
@@
Optuna tries to maximize forget set loss (or WMDP accuracy) after relearning with supervised fine-tuning. While we are interested in the overall recoverability of unlearned behavior (e.g., through jailbreaks or out-of-distribution attacks), here we focus on supervised fine-tuning, as it is the most straightforward and most reliable way to resurface removed behavior (Lynch et al., 2024).
@@
For the WMDP experiments, rather than controlling for the retain loss increase, we use the accuracy on the MMLU benchmark (Hendrycks et al., 2021) to gain a comprehensive understanding of performance preservation.  [^2] We use temperature=1 to get MMLU answers, to avoid discontinuous answer switching which happens with temperature=0, which makes trial rejection less predictable. During unlearning we allow at most 1 percentage point drop in MMLU accuracy.
@@
Figure 4: Accuracy on WMDP-Bio after unlearning and relearning on Pile-Bio. The base level on the right (45.5%) is the accuracy with no unlearning applied, but after relearning. Reported accuracy is the average of the last 60 valid trials in each hyperparameter search. We only show Llama, because smaller models had near-random accuracy on WMDP-Bio. Here, we use MUDMAN with negative entropy (Tamirisa et al., 2024a) as the unlearning loss. Similarly to Figure 3, using Disruption Masking is crucial (orange vs red) and it accounts for most of the improvement over TAR. There is also clear improvements from meta-learning and unlearning gradient normalization.
@@ 4 Building a robust unlearning pipeline
Among them, we identify several key components that consistently improve unlearning robustness. We also propose a novel method called Disruption Masking. We integrate these into MUDMAN (Meta-Unlearning with Disruption Masking and Normalization), which outperforms the state-of-the-art TAR method Tamirisa et al. (2024a). See Algorithm 1 for pseudocode and Appendix F for a PyTorch implementation. Below, we describe each component of MUDMAN and its overall results.
@@ 4.1 Meta-learning is effective
First, we confirm the effectiveness of MAML (model-agnostic meta-learning) (Finn et al., 2017), which is increasingly used in modern unlearning methods (Tamirisa et al., 2024a; Henderson et al., 2023; Tamirisa et al., 2024b). MAML involves training a copy of the main model, which we call the adversary, on the forget set and applying its gradients to the main model.
Despite its adoption, no mechanistic explanation for why MAML helps with robust unlearning has been proposed. We hypothesize that while naive unlearning merely makes harmful circuits dormant (Lee et al., 2024) (and dormant circuits cannot be further unlearned because backpropagation cannot locate them), MAML trains an adversary in which these dormant circuits are reactivated, so that backpropagation can continue to erase them.
Traditional MAML employs an inner loop to train multiple adversaries and accumulate unlearning gradients before updating the main model (Finn et al., 2017). However, in early experiments, we found that interleaving adversary and main model updates actually improves performance. Therefore, we flatten the process into a single loop (see Algorithm 1) and focus on training a single adversary deeply, rather than multiple ones briefly.
@@ 4.2 Disruption Masking drives performance
This is our key contribution from MUDMAN. In existing unlearning techniques (Tamirisa et al., 2024a; Rosati et al., 2024; Li et al., 2024), the model is also trained on the retain set during unlearning, hoping to revert any unintended disruptions. Instead, we adopt a more selective approach, aiming to avoid disruptions altogether rather than correcting them later.
@@
Implementation-wise, to generalize beyond the current batch and identify what might break retain set performance overall, we use SGD momentum (Rumelhart et al., 1986), which adds retain gradients to a decaying accumulator. Then, instead of the signs of the single-batch retain gradients, we use the signs of this accumulator. This incurs no additional memory cost, as we already need to store retain gradients alongside unlearning gradients, so we simply store the accumulator instead.
@@ 4.5 Improvements from MUDMAN
We validate the superiority of MUDMAN on three models and two unlearning tasks over TAR (Tamirisa et al., 2024a), a state-of-the-art unlearning method. For a clean comparison, we adapt the TAR implementation to match our setup: using a single training loop and omitting its initial representation noising step. As shown in Figure 3, MUDMAN consistently outperforms this adapted TAR across all tested cases. In individual setups, while some ablations occasionally match the full MUDMAN, none of them ever significantly surpasses it.
@@ 6 Limitations
In this study, we only use supervised fine-tuning to elicit unwanted behavior. While it is the most powerful and reliable elicitation method (Lynch et al., 2024), future work could extend to evaluate other attacks, in which the attacker does not have weight access, for example with jailbreak prompts (Zou et al., 2023).
To further test whether our techniques truly remove capabilities or just make them harder to recover, one could use the approach introduced by Deeb and Roger (2024), where the attacker tries to uncover some unknown, supposedly unlearned facts, by fine-tuning on another, non-overlapping set of facts.
@@
For simplicity, we only evaluate monolithic methods which apply a single algorithm throughout the whole unlearning process. This involved stripping our TAR baseline (Tamirisa et al., 2024a) of its initial representation noising step and only keeping the meta-learning core of the algorithm. Future work could investigate the effectiveness of consecutively stacking different unlearning methods, looking for synergies between them.
@@
Filip Sondej’s work was partially funded by a grant from Open Philanthropy.
@@ Appendix A Future work
Finally, the concept of unlearning could be inverted, and we could try to make sure that some critical desired behavior always happens, and will never be disturbed out-of-distribution or under jailbreaks or fine-tuning. This is particularly useful in the deliberative alignment framework [Guan et al., 2025], where we want some lines of reasoning to trigger reliably in certain contexts.
@@ Appendix D Unlearning tendencies
We had a partial success unlearning dangerous tendencies. We chose cruelty as our candidate, and as the forget set we used animal abuse examples from BeaverTails [Ji et al., 2023]. It contains Q&A pairs, with questions like "What are some ways to torture a small animal?". As the retain set, we used pairs from BeaverTails with benign answers.
@@ E.1 Dampening relearning gradients
Relearning will not happen if the gradients during relearning are near zero. We tried several techniques to dampen them, which we describe below. Unfortunately they were outperformed by more straightforward techniques just relying on backpropagation. This failure is actually consistent with findings by Finn et al. [2017], who have shown that meta-learning which uses second-order derivatives does not perform any better than the simpler and cheaper first-order method (model-agnostic meta-learning).
@@ E.2 Direct weight edits
We can locate which MLP neurons are most active on the forget task, and at the same time least active on the retain task, and then just ablate them. This is similar to existing model editing methods [Wu et al., 2023, Suau et al., 2024].
@@ E.4 Making unlearning more selective
To make calculation of disruption more accurate, we can also augment the normal retain loss with a loss aiming to leave the model activations unchanged Tamirisa et al. [2024a], Zou et al. [2024]. In our setup we saw no improvement when using this activation loss, but maybe some effect will become visible after we also look at the performance on capability benchmarks.
@@
It is possible to draw some parallels between Disruption Masking and A-GEM technique [Chaudhry et al., 2019] from the continual learning field. Similarly, the method in the continual learning field that corresponds to this aggregation of absolute values, would be Online-EWC [Schwarz et al., 2018], which tries to estimate disruption on previous tasks by squaring the gradients.

### v5

# AI That Keeps Up: Workshop on Continual and Compatible Foundation Model Updates (CCFM)

@@ Abstract
Additionally, we identify the need for normalizing the unlearning gradients, and also confirm the usefulness of meta-learning. We combine these insights into MUDMAN (Meta-Unlearning with Disruption Masking and Normalization) and validate its effectiveness at preventing the recovery of dangerous capabilities. MUDMAN outperforms the prior TAR method by 40%, setting a new state-of-the-art for robust unlearning. Code: github.com/filyp/MUDMAN
@@ 1 Introduction
Language models can acquire dangerous skills during pre-training, such as manipulation, hacking and even knowledge useful for creating bioweapons (li_wmdp_2024). They may also learn about the safeguards used to control them, which in the future may enable them to subvert such safeguards (greenblatt_alignment_2024; roger_case_2024).
Studies show that popular safety fine-tuning techniques like DPO and RLHF fail to remove dangerous knowledge; instead, they minimally modify model weights to hide it (lee_mechanistic_2024), allowing it to reemerge through jailbreak inputs (zou_universal_2023), or even accidentally (qi_fine-tuning_2023; deeb_unlearning_2024). Even specialized unlearning techniques turn out to be easily reversible (lynch_eight_2024; lucki_adversarial_2025; deeb_unlearning_2024). It remains unclear which components of unlearning algorithms can make harmful capabilities truly irreversible.
@@ 2 Related work
Some unlearning approaches disrupt intermediate activations within the model (zou_improving_2024; li_wmdp_2024; rosati_representation_2024), while others attempt to locate and ablate weights responsible for unwanted behavior wang_large_2024; wu_depn_2023; uppaal_detox_2024; suau_whispering_2024. However, lo_large_2024 found that even when unwanted concepts are directly removed, the model can quickly learn to represent them again using neurons with similar meaning.
To address this, modern unlearning techniques increasingly incorporate MAML (model-agnostic meta-learning) (tamirisa_tamper-resistant_2024; henderson_self-destructing_2023; tamirisa_toward_2024). This approach anticipates how an attacker could relearn the target capability, by deriving unlearning gradients from a copy of the model trained on the forget set (finn_model-agnostic_2017).
However, for each existing unlearning methods, there are still ways to elicit the supposedly removed capabilities, such as jailbreak prompts, few-shot prompting, fine-tuning, in-context learning, out-of-distribution inputs or disabling refusal mechanisms with representation engineering (lynch_eight_2024; lucki_adversarial_2025).
@@ 3 Experiment setup
We conduct experiments with three language models of different sizes: pythia-14m (biderman_pythia_2023). SmolLM-135M (allal2024SmolLM) and Llama-3.2-1B (grattafiori2024llama). Using pythia-14m allowed us to iterate fast on our techniques, then we validate our findings on the larger, more modern models.
@@
We demonstrate skill unlearning by attempting to unlearn Python coding ability by using function examples from CodeSearchNet (husain2019codesearchnet), while retaining the performance on Wikitext (merity2016pointer). To ensure we do not unintentionally unlearn English, we remove comments and docstrings from the Python dataset. While programming is not a harmful behavior, using loss on the Python dataset provides a high signal-to-noise ratio for comparing unlearning methods and serves as a good baseline.
Then we move to knowledge unlearning with a realistic harmful target: unlearning biohazardous knowledge using Pile-Bio (tamirisa_tamper-resistant_2024), a subset of the Pile (gao_pile_2020) containing texts about molecular biology. The rest of the Pile serves as the retain set.
To test unlearning effectiveness on Pile-Bio, we measure the accuracy on WMDP-Bio (li_wmdp_2024), a dataset of 1273 multiple-choice questions designed as a proxy measurement of hazardous biosecurity knowledge. We use temperature=0 to generate the answers.
@@
Method performance is highly dependent on hyperparameter selection. To ensure a fair comparison, we perform an automated hyperparameter search for each tested method using Optuna (akiba_optuna_2019). We verify that each search converges and that the chosen hyperparameter ranges are not saturated. Each search includes hundreds of trials, with each trial consisting of an unlearning stage followed by a fixed relearning stage, and the amount of compute in each stage is constant. (See Appendix G for details.)
@@
Optuna tries to maximize forget set loss (or WMDP accuracy) after relearning with supervised fine-tuning. While we are interested in the overall recoverability of unlearned behavior (e.g., through jailbreaks or out-of-distribution attacks), here we focus on supervised fine-tuning, as it is the most straightforward and most reliable way to resurface removed behavior (lynch_eight_2024).
@@
For the WMDP experiments, rather than controlling for the retain loss increase, we use the accuracy on the MMLU benchmark (hendrycks_measuring_2021) to gain a comprehensive understanding of performance preservation.  [^2] We use temperature=1 to get MMLU answers, to avoid discontinuous answer switching which happens with temperature=0, which makes trial rejection less predictable. During unlearning we allow at most 1 percentage point drop in MMLU accuracy.
@@
Figure 4: Accuracy on WMDP-Bio after unlearning and relearning on Pile-Bio. The base level on the right (45.5%) is the accuracy with no unlearning applied, but after relearning. Reported accuracy is the average of the last 60 valid trials in each hyperparameter search. We only show Llama, because smaller models had near-random accuracy on WMDP-Bio. Here, we use MUDMAN with negative entropy (tamirisa_tamper-resistant_2024) as the unlearning loss. Similarly to Figure 3, using Disruption Masking is crucial (orange vs red) and it accounts for most of the improvement over TAR. There is also clear improvements from meta-learning and unlearning gradient normalization.
@@ 4 Building a robust unlearning pipeline
Among them, we identify several key components that consistently improve unlearning robustness. We also propose a novel method called Disruption Masking. We integrate these into MUDMAN (Meta-Unlearning with Disruption Masking and Normalization), which outperforms the state-of-the-art TAR method tamirisa_tamper-resistant_2024. See Algorithm 1 for pseudocode and Appendix F for a PyTorch implementation. Below, we describe each component of MUDMAN and its overall results.
@@ 4.1 Meta-learning is effective
First, we confirm the effectiveness of MAML (model-agnostic meta-learning) (finn_model-agnostic_2017), which is increasingly used in modern unlearning methods (tamirisa_tamper-resistant_2024; henderson_self-destructing_2023; tamirisa_toward_2024). MAML involves training a copy of the main model, which we call the adversary, on the forget set and applying its gradients to the main model.
Despite its adoption, no mechanistic explanation for why MAML helps with robust unlearning has been proposed. We hypothesize that while naive unlearning merely makes harmful circuits dormant (lee_mechanistic_2024) (and dormant circuits cannot be further unlearned because backpropagation cannot locate them), MAML trains an adversary in which these dormant circuits are reactivated, so that backpropagation can continue to erase them.
Traditional MAML employs an inner loop to train multiple adversaries and accumulate unlearning gradients before updating the main model (finn_model-agnostic_2017). However, in early experiments, we found that interleaving adversary and main model updates actually improves performance. Therefore, we flatten the process into a single loop (see Algorithm 1) and focus on training a single adversary deeply, rather than multiple ones briefly.
@@ 4.2 Disruption Masking drives performance
This is our key contribution from MUDMAN. In existing unlearning techniques (tamirisa_tamper-resistant_2024; rosati_representation_2024; li_wmdp_2024), the model is also trained on the retain set during unlearning, hoping to revert any unintended disruptions. Instead, we adopt a more selective approach, aiming to avoid disruptions altogether rather than correcting them later.
@@
Implementation-wise, to generalize beyond the current batch and identify what might break retain set performance overall, we use SGD momentum (rumelhart_learning_1986), which adds retain gradients to a decaying accumulator. Then, instead of the signs of the single-batch retain gradients, we use the signs of this accumulator. This incurs no additional memory cost, as we already need to store retain gradients alongside unlearning gradients, so we simply store the accumulator instead.
@@ 4.5 Improvements from MUDMAN
We validate the superiority of MUDMAN on three models and two unlearning tasks over TAR (tamirisa_tamper-resistant_2024), a state-of-the-art unlearning method. For a clean comparison, we adapt the TAR implementation to match our setup: using a single training loop and omitting its initial representation noising step. As shown in Figure 3, MUDMAN consistently outperforms this adapted TAR across all tested cases. In individual setups, while some ablations occasionally match the full MUDMAN, none of them ever significantly surpasses it.
@@ 6 Limitations
In this study, we only use supervised fine-tuning to elicit unwanted behavior. While it is the most powerful and reliable elicitation method (lynch_eight_2024), future work could extend to evaluate other attacks, in which the attacker does not have weight access, for example with jailbreak prompts (zou_universal_2023).
To further test whether our techniques truly remove capabilities or just make them harder to recover, one could use the approach introduced by deeb_unlearning_2024, where the attacker tries to uncover some unknown, supposedly unlearned facts, by fine-tuning on another, non-overlapping set of facts.
@@
For simplicity, we only evaluate monolithic methods which apply a single algorithm throughout the whole unlearning process. This involved stripping our TAR baseline (tamirisa_tamper-resistant_2024) of its initial representation noising step and only keeping the meta-learning core of the algorithm. Future work could investigate the effectiveness of consecutively stacking different unlearning methods, looking for synergies between them.
@@
Filip Sondej’s work was partially funded by a grant from Coefficient Giving.
@@ Appendix A Future work
Finally, the concept of unlearning could be inverted, and we could try to make sure that some critical desired behavior always happens, and will never be disturbed out-of-distribution or under jailbreaks or fine-tuning. This is particularly useful in the deliberative alignment framework [guan_deliberative_2025], where we want some lines of reasoning to trigger reliably in certain contexts.
@@ Appendix D Unlearning tendencies
We had a partial success unlearning dangerous tendencies. We chose cruelty as our candidate, and as the forget set we used animal abuse examples from BeaverTails [ji_beavertails_2023]. It contains Q&A pairs, with questions like "What are some ways to torture a small animal?". As the retain set, we used pairs from BeaverTails with benign answers.
@@ E.1 Dampening relearning gradients
Relearning will not happen if the gradients during relearning are near zero. We tried several techniques to dampen them, which we describe below. Unfortunately they were outperformed by more straightforward techniques just relying on backpropagation. This failure is actually consistent with findings by finn_model-agnostic_2017, who have shown that meta-learning which uses second-order derivatives does not perform any better than the simpler and cheaper first-order method (model-agnostic meta-learning).
@@ E.2 Direct weight edits
We can locate which MLP neurons are most active on the forget task, and at the same time least active on the retain task, and then just ablate them. This is similar to existing model editing methods [wu_depn_2023, suau_whispering_2024].
@@ E.4 Making unlearning more selective
To make calculation of disruption more accurate, we can also augment the normal retain loss with a loss aiming to leave the model activations unchanged tamirisa_tamper-resistant_2024, zou_improving_2024. In our setup we saw no improvement when using this activation loss, but maybe some effect will become visible after we also look at the performance on capability benchmarks.
@@
It is possible to draw some parallels between Disruption Masking and A-GEM technique [chaudhry_efficient_2019] from the continual learning field. Similarly, the method in the continual learning field that corresponds to this aggregation of absolute values, would be Online-EWC [schwarz_progress_2018], which tries to estimate disruption on previous tasks by squaring the gradients.